### v1

# Exploiting Student Parallelism for Low-latency GPU Inference of BERT-like Models in Online Services

Abstract.
Due to high accuracy, BERT-like models have been widely adopted by discriminative text mining and web searching. However, large BERT-like models suffer from inefficient online inference, as they face the following two problems on GPUs. First, they rely on the large model depth to achieve high accuracy, which linearly increases the sequential computation on GPUs. Second, stochastic and dynamic online workloads cause extra costs.
[…]
1. Introduction
Due to the marvelous success in natural language understanding, BERT-like models (devlin2018bert, ; liu2019roberta, ; he2021deberta, ) have been the foundation model for discriminative text mining and web searching in various online services. However, large BERT-like models pay the price of high inference latency for the model capability. Besides high accuracy, a wide range of web services are time-sensitive and require efficient online inference. For example, news sentiment classification for automatic stock trading should be quicker than any others, because even a sub-millisecond delay can determine the profits (aquilina2022quantifying, ); web searching usually has a ten-millisecond-level time budget for every query to score and rank the relevance of document candidates (yang2020model, ); and online writing assistants need to give immediate response to text streams while users are typing (DBLP:conf/acl/KanekoMKSI20, ).
Some existing works focus on reducing the computation complexity of BERT-like models such as the compact model design (sun2020mobilebert, ; xu2021bert, ), knowledge distillation (sanh2019distilbert, ; sun2019patient, ; jiao2019tinybert, ), and pruning (DBLP:conf/rep4nlp/GordonDA20, ; michel2019sixteen, ; hou2020dynabert, ) which effectively reduce the model size for weak edge devices. However, since deep learning essentially relies on model depth for high model capability (lecun2015deep, ), these methods prefer reducing model width to model depth (about half the layers remained) to better maintain the accuracy. Moreover, early exiting methods (DBLP:conf/acl/XinTLYL20, ; DBLP:conf/acl/LiuZWZDJ20, ; zhou2020bert, ) still have some hard samples to go through all layers. Therefore, their deep architectures largely degrade their inference performance on GPUs which are the de-facto platform for deep learning models. Figure 1(a) shows that model depth linearly increases the latency due to serial forward propagation, while GPU has enough parallel computation capability to keep the same latency and throughput for different model widths as shown in Figure 1(b). Therefore, these deep and thin models are not suitable for efficient GPU inference.
Furthermore, previous works pay extra costs for the stochastic online workloads. Different from collected data batches in training, online inference samples of different lengths arrive in stochastic and dynamic patterns. Since there are occasional workload bursts due to breaking events, previous works have to reserve over-provisioned GPUs idle for the most time (zhang2019mark, ; DBLP:conf/usenix/Romero0YK21, ; 276950, ). Additionally, Figure 1(c) indicates that GPUs need large batches for high throughput, because latency remains stable and throughput keeps increasing until the batch size is large enough. For high throughput and utilization, existing works employ a waiting queue to batch enough data samples from stochastic online workloads (zhang2019mark, ; ali2020batch, ; DBLP:conf/usenix/Cui00WLZ0G22, ), so early-arrived samples have higher latency due to the waiting. And text sequences have different lengths ranging from 10 to hundreds, but data parallelism on GPUs requires all samples in one batch to have the regular length. Existing works commonly pad short samples into the maximum length to form the regular batch (devlin2018bert, ). As shown in Figure 1(d), latency can increase rapidly with sequence length due to $O(N^{2})$ complexity of self-attention (DBLP:conf/nips/ZaheerGDAAOPRWY20, ), so paddings also bring in significant extra computations.
(a) Model Depth (Layer #)
(b) Model Width (Hidden Dim.)
(c) Input Batch Size
(d) Input Sequence Length
Figure 1. Influences of different model and input factors on latency (red solid line) and throughput (green dashed line)
Therefore, we present Student Parallelism to explore student parallelism for low-latency GPU inference of BERT-like models, especially for the web services having stochastic online workloads. At the core of Student Parallelism, it distills the original deep model into a parallel group of shallow student models. Therefore, the sequential layer computations can be replaced with the equivalent parallel computations of students without compromising accuracy. To achieve this, we propose a novel knowledge distillation method based on explicit boosting ensemble and stacking distillation (§ 3.2). Despite the parallel inference, every student is sequentially trained to gradually reduce the error between the teacher and the group of students. Furthermore, every student distills all previous students into its intermediate layer by stacking distillation, so that all students can make better corrections with the estimated output of previous students. Therefore, all students can maintain the accuracy with no more than one quarter of the original layers (even only two layers in some cases), which significantly reduces the inference latency on GPUs. Additionally, it conducts adaptive student pruning to improve the generalization of the remained students and drop some last-added students with minimal accuracy loss, which can temporarily improve throughput for the occasional workload burst.
To reduce the extra costs of batching from online workloads, Student Parallelism employs specialized system designs for student parallelism (§ 3.3). It allocates the students of the same group on different GPUs to better parallelize even the single sample. It further replicates the student group into multiple copies to run concurrently and share the same set of GPUs with MPS. Therefore, it enables immediate and concurrent inference of different samples without batching and interference to minimize waiting and paddings.
[…]
Figure 2. Comparison in the trade-off between accuracy and latency (the upper left corner is optimal)
In summary, our work makes the following contributions:
  1. (1)
    We propose student parallelism to distill the deep model into a group of parallel students with lower depth than baselines (e.g., two layers) for low latency with comparable accuracy.
  2. (2)
    To handle online workloads, we can reduce the student number with minimal accuracy loss for workload burst, and we also make specialized system designs for student parallelism to reduce waiting and paddings.
  3. (3)
    We conduct comprehensive experiments to verify the effectiveness of our Student Parallelism, where it outperforms other baselines in terms of accuracy, latency, and throughput.
2. Related Work
Previous works have explored various efficient inference methods for BERT-like models. However, they still suffer from high latency due to large model depth and extra costs for online workloads. There are three categories of existing works as follows:
  • Model Compression: these works only generate a single small model for weak edge devices, which still suffer from considerable model depth and high latency on GPUs. MobileBERT (sun2020mobilebert, ) and NAS-BERT (xu2021bert, ) make compact model design that can achieve the similar accuracy. Knowledge distillation transfers the knowledge from a large teacher model to a small student model (sanh2019distilbert, ; sun2019patient, ; jiao2019tinybert, ), so that the small student mimics the behavior of its teacher. Pruning can remove redundant weights, neurons, attention heads or tokens in well-trained models for less costs and better generalization (hou2020dynabert, ; malach2020proving, ; kim2022learned, ). Early exiting (DBLP:conf/acl/XinTLYL20, ; DBLP:conf/acl/LiuZWZDJ20, ; zhou2020bert, ) adds multiple classifiers on different layers as the multiple exits, so that it allows early exit from bottom layers for ”easy” samples. Quantization (kim2021bert, ; shen2020q, ; zafrir2019q8bert, ; tang2022mkq, ) replaces slow float operators with fast integer operators, which releases higher parallel computation capability. However, deep learning essentially relies more on model depth than model width for high model capability (lecun2015deep, ), so they still require a considerable large model depth to maintain the accuracy. For example, MobileBERT and NAS-BERT are still deep (12 or even 24 layers) but much thinner to avoid accuracy loss. 6-layer DistillBERT can maintain 96.3% overall score of the 12-layer Bert teacher, but the 3-layer PKD-BERT can only get 88.5% (§ 4.2.1). Model pruning (hou2020dynabert, ) can only reduce about the half of layers at most to maintain good accuracy. In early exiting, the ”difficult” samples still have to go through all layers. Quantization not only needs hardware support in low-bit integers (zafrir2019q8bert, ) but also fails to reduce the model depth and computation complexity. It also relies on large batch sizes to increase parallelism degree and speed-up (e.g., the speedup is 2.4×2.4$\times$2.4 × and $\times3$ when the batch size is 1 and 8 respectively (kim2021bert, )).
  • Weak Model Ensemble: Clipper (DBLP:conf/nsdi/CrankshawWZFGS17, ) and Cocktail (276950, ) adopt bagging ensemble of different relatively small models to run in parallel for lower latency and high accuracy, which can only select the models readily trained in the above-mentioned methods. However, they rely on the different model architectures for the ensemble diversity, resulting in an imbalanced workload and straggler problem. The Mixture-of-Experts (MoE) in attention heads (DBLP:conf/iclr/ShazeerMMDLHD17, ; DBLP:conf/naacl/ZuoZLHZC22, ) has a router to activate only one or two experts in every layer. It can keep constant inference cost no matter how wide the model is, but it cannot decrease the model depth. SenseAI (wang2021sensai, ) divides the multi-class model into several pruned binary models, which still have a large model depth and irregular pruned architectures. It cannot be used in regression and binary classification tasks either. Some collaborative and self-distillation (DBLP:conf/icann/ZainJZ22, ; sun2021collaborative, ) explore distilling multiple teachers and existing students into a new better student for the time-shifting distribution, but they do not leverage multiple parallel students together for lower model depth.
  • System Designs for Online Workloads: existing system works pay extra costs to handle dynamic online workloads. They reserve some over-provisioned GPUs in advance for potential workload bursts, which are underutilized for the most time. According to workload prediction and profiling, previous works can dynamically set the batch size and maximum waiting time to reduce the idle waiting (zhang2019mark, ; ali2020batch, ), but it can hardly be optimal. Some recent works (triton, ; pytorch2018pytorch, ; abadi2016tensorflow, ; zhai2022bytetransformer, ) bring in some operators supporting ragged batches to avoid extra paddings, but they can hardly support and perfectly optimize all operators and situations (DBLP:conf/mlsys/FegadeCGM22, ).
In contrast, Student Parallelism trains a group of parallel students for GPUs other than a single student for weak edge devices, which can achieve the lowest model depth. It can even effectively reduce the model depth of large models like BERT-large and DeBERTa-xlarge, while the others fail. All students also have the same shallow architectures to be free of stragglers. Without over-provisioned GPUs, it can reduce the student number for higher throughput in case of workload bursts. Due to student parallelism, it also saves the costs of waiting queues and padding. Being orthogonal to quantization, it can quantize all students for further improvement.
3. Methodology
In this section, we first introduce the framework of Student Parallelism. Then we discuss how to offline train the equivalent group of parallel students for student parallelism. Finally, we describe how the multiple students run in parallel for online inference.
3.1. Framework Overview
We propose Student Parallelism to distill the original teacher model into a group of shallow student models for student parallelism, instead of a single student model. The core idea of Student Parallelism is to trade more parallel student models for less sequential layers to maintain accuracy. Specifically, Student Parallelism has the following two different phases to realize different goals:
  • Offline Training (§ 3.2): It aims to maintain the accuracy with the adaptive numbers of parallel students in low model depth. To be friendly to GPU acceleration, every student share the same shallow model architecture, resulting in weak model capability and homogeneity. To improve model capability and diversities, all students are sequentially trained on different sub-sampled subsets and virtually stacked via stacking distillation to gradually reduce the residual errors. By dynamically dropping some last-trained students, Student Parallelism conducts adaptive student pruning to improve the generalization of the remained students and minimize the accuracy loss. Then it can temporally reduce the student number to handle workload burst.
  • Online Inference (§ 3.3): it distributes the students of the same groups on the multiple GPUs to run in parallel for the single sample. Student Parallelism properly allocates the student group on GPUs and efficiently processes inference samples from online workload. And it replicates the student group into multiple copies to share GPUs with the support of MPS, so that Student Parallelism enables the immediate and concurrent inference of different individual samples without large data batches. If necessary, it employs the length-aware buffer to merge massive concurrent samples in similar lengths for higher throughput.
3.2. Offline Training
There are two challenges in training the equivalent group of parallel and shallow students for efficient GPU inference without affecting accuracy. One is how to maintain accuracy with those weak and homogeneous students having much fewer layers for low latency. The other is how to reduce student numbers with minimal accuracy loss for temporary high throughput. To overcome the two challenges, Student Parallelism employs the following two-stage training:
  1. (1)
    Sequential student training (line: 1-14): To generate a similar final pooled representation, all students are sequentially trained to gradually reduce the error between the teacher and all the students in boosting ensemble. To improve ensemble diversity and accuracy, it proposes different training losses, model initialization, and training datasets. All students are virtually stacked by the stacking distillation to improve the results.
  2. (2)
    Adaptive student pruning (line: 15-29): To temporarily improve throughput with minimal accuracy loss, Student Parallelism proposes adaptive student pruning to reduce the student number. By following the same way of sub-network training in NAS, it drops different numbers of last-added students to extract representations during the training. For every training data batch, it trains the output classifier upon different numbers of students to mimic the teacher’s prediction behavior.
3.2.1. Sequential Training of Student Models
Figure 3. The sequential training of virtually stacked students to gradually reduce the residual errors on the final representation between the teacher and all students.
Now we introduce how Student Parallelism trains the multiple parallel and shallow students to extract similar final representations with the original deep teacher. According to ensemble learning theory, different weak models need significant diversities to improve generalization and accuracy (zhou2021ensemble, ). Since all student models share the same shallow network for student parallelism, we propose the training strategy as described in Algorithm 1 to improve diversities from other aspects, including training loss, parameter initializations, and data subsets.
To achieve diversity, every student has different training loss functions applied to the final representation and intermediate representation (line 9) , as illustrate on Figure 3. All students are sequentially trained in the style of gradient boosting ensemble to gradually correct its own residual error on the final representation. Except that the first student directly mimics the final representation of the teacher, the $i-th$ student learns to correct the error of the ensemble of its own previous students. We keep adding new students until the overfitting happens according to the validation on spared test data (line 3). Formally, we define the boosting ensemble of the $M$ students having $N$ attention layers to generate the final representation as follows:
(1) $B^{(M)}(x):=\sum_{m=0}^{M}\alpha^{(m)}S^{(m)}_{N}(x)$
where $S^{(m)}_{N}$ is the final representation of m-th student model in its $N$-th layer, and $\alpha^{(m)}$ is its multiplier . Then we can optimize the following residual training loss $L_{\text{boost}}^{(i)}$ for the i-th student :
(2) $L_{\text{boost}}^{(i)}(x)=\frac{1}{2}\left\|T(x)-B^{(i-1)}(x)-S^{(i)}_{N}(x)\right\|^{2}.$
In Equation 2, the $T(x)$ stands for the final representation output of the teacher model and $B^{(i-1)}(x)$ is the boosting ensemble of its all previous $i-1$ students. And only the i-th training student $S^{(i)}$ is trainable, while the other models like $T$ and $B^{(i-1)}(x)$ are fixed. After training the student $S^{(i)}$, we can set its multiplier $\alpha_{i}$ by line search to minimize the loss $\frac{1}{2}(T(x)-B^{(i)}(x))^{2}$, except that $\alpha_{0}$ is always 1. And we make the mathematical interpretation and formal proof on the better convergence of the final representations of all students based on the gradient boosting ensemble in the appendix.
Algorithm 1 Sequential Student Training
1:TeacherModel, PreMiniModel
2:#Sequential training of student models
3:studentList = [PreMiniModel]
4:while not overfitting do
5:     newStudent = studentList[-1]
6:     Sample the new training dataset
7:     while training do
8:         outputs_t = TeacherModel(x)
9:         outputs_s = newStudent(x)
10:         Compute the losses as Equation 2 and Equation 3
11:         Update the newStudent
12:     end while
13:     studentList.append(copy(newStudent))
14:end while
15:studentList = studentList[:-1]
Despite that all students run independently, we propose stacking distillation to make every student virtually stacked on its own previous students. Any student learns to distill the ensemble of its all previous students into the representation of its $\lceil N/2\rceil$-th layer. Since stacking distillation makes the intermediate representation imitates all the previous students, the following layers can better correct the error by refining the intermediate representation, in the similar way of a deep BERT-like model. The loss $L_{\text{stack}}^{(i)}$ of stacking distillation for the i-th student is defined as follows:
(3) $L_{\text{stack}}^{(i)}(x)=\frac{1}{2}\left\|B^{(i-1)}(x)-S^{(i)}_{\lceil N/2\rceil}(x)\right\|^{2}$
where $S^{(i)}_{\lceil N/2\rceil}(x)$ denotes the intermediate representation of the $N/2$-th layer in the i-th student model. In Equation 3, only $S^{(i)}$ is trainable while the other models are fixed.
Therefore, the overall training loss $L$ for the whole representation distillation can be written as follows:
(4) $L=L_{\text{boost}}+\lambda L_{\text{stack}}$
where $\lambda$ is the hyper-parameter to balance the final representation distillation and stacking distillation.
Besides the different training loss, Student Parallelism also brings in diversities from different perspectives. All students have different initialized parameters. The first student is initialized with a pre-trained language model with limited layers to improve the generalization, while any other student copies the parameters from the last trained student for the initialization (line 4). Additionally, every student model is trained on different sampled subsets of training data (line 5), except only the first student is trained on the whole training data. For any student, we pick up the top $a\%$ samples with the largest residual errors between the teacher and the existing students. And we randomly sample another $b\%$ of all the training data. And the subset of training data is augmented in the way suggested by previous works (jiao2019tinybert, ; hou2020dynabert, ).
3.2.2. Adaptive Pruning of Student Models
Algorithm 2 Adaptive Students Pruning
1:#Adaptive prunning of student models
2:while training do
3:     softLabel = TeacherModel(batch)
4:     for $k\leftarrow 1$ to StudentNum do
5:         rep=$\sum_{j}^{k}\alpha_{j}studentList[j](batch)$
6:         output = Classifer(rep)
7:         loss = softCrossEntropy(output,softLabel)
8:         Accumulate the gradients of current k students
9:     end for
10:     Update all students based on the accumulated gradients
11:end while
12:for $k\leftarrow 1$ to StudentNum do
13:     evaluate the previous $k$ students on validation set
14:end for
15:Choose the best previous $n$ students
To improve throughput with minimal accuracy loss, Student Parallelism further employs adaptive student pruning to reduce the student number and improve the generalization of the remained students. We follow the sub-network training in Network Architecture Searching (NAS) (cai2019once, ) and DynaBERT (hou2020dynabert, ) to prune the redundant students in boosting ensemble for final representation extraction. Then it trains the output classifier on the final representations extracted from different numbers of students to mimic the prediction output of the teacher. Since the newly added students of boosting ensemble gradually improve accuracy by reducing the residual error, we drop and prune last-trained students to reduce overall computation for higher throughput with minimal accuracy loss. Moreover, pruning can even improve accuracy and generalization by only keeping the suitable number of students, because Occam’s razor recommends searching for hypotheses constructed with the smallest possible set of elements. In § 6, we verify the reduced accuracy loss of different numbers of students and better generalization of the best-pruned set of students.
Specifically, we state the details about adaptive student pruning in Algorithm 2. During the training of one data batch, It dynamically samples different numbers of students to train simultaneously (line 4). In this stage, different numbers of ensembled students generate the final representation (line 5), which has quite limited communication costs thanks to the small size of the pooled final representation. Then we put a new classifier layer on the sum of the selected students’ outputs (line 6). And we take the soft labels generated by the teacher model (line 3) as the target in the soft cross-entropy loss rather than the hard labels of the ground truth (line 7). Since there are multiple forward and back propagations of different numbers of students for the same training batch, it accumulates all the gradients to update the parameters of the classifier and all students (line 8 and 10). Finally, we evaluate different numbers of previous $k$ students on the validation set to choose the optimal student number for the best pruned students, instead of keeping all students (line 12-15).
3.2.3. Training Cost Discussion
Our offline training phase has a similar time cost to other model compressions. Although we need to train multiple students, the student is a much smaller pre-trained model compared with other baselines. And each student is only trained on the sub-sampled dataset based on the residual error in the boost ensemble learning. We can also use distributed training on multiple GPUs to speed up. Generally, Student Parallelism has a similar level of offline training cost with previous distillation works (about the 1.35x training time of DynaBERT). And the training cost can be amortized during the long-term deployment.
3.3. Online Inference
Figure 4. The online inference system designs for student parallelism: model allocation of parallel student groups and inference sample dispatching.
For efficient GPU inference for stochastic online workloads. we make specialized system designs to realize student parallelism, including student allocation and sample dispatching. We assume that the GPU cluster running online inference has multiple nodes, and each node has multiple GPUs. In general, we have one global dispatcher to balance the workloads among all nodes. And every node runs its own resource management and task scheduling. As shown in Figure 4, Student Parallelism has the following procedures during the online inference phase:
  1. (1)
    In the initialization, the resource manager starts multiple student groups running in parallel, and every group distributes its students over multiple GPUs.
  2. (2)
    User submits a request triggering one or more inference samples.
  3. (3)
    The dispatcher assigns the samples to different GPU nodes in the round-robin manner.
  4. (4)
    After reaching the GPU node, the sample is pre-processed and then enters the length-aware buffer. And if the buffer is not empty, it tries to merge the samples of the similar length into small batches.
  5. (5)
    Whenever any student group is idle, it immediately sends out the first buffered data sample or batch of samples in similar lengths for immediate inference without waiting.
  6. (6)
    According to the dynamic intensity of online workload, resource management decreases the student number for larger throughput or increases it for better accuracy.
  7. (7)
    It returns results to users.
3.3.1. Student Allocation
Figure 5. Batch VS. Concurrent Samples: the top is the timeline of batching and processing all samples without overlapping, and the bottom is how multiple models concurrently process different samples without batching to overlap the data transferring and inference computations.
Resource management employs a specialized student allocation strategy for the group of parallel students in every node. Since all student models share parallel and homogeneous network architecture, it can distribute the different students in the same group over the multiple GPUs in the same node as shown in Figure 4, which can better parallelize the inference of every sample. Although the parallel students in the same group need communication to gather the final representation, it only takes about 0.2 ms due to the small size of the final pooled representation and fast NV-link or PCI-E.
Student Parallelism further replicates the student group into multiple copies and assign the multiple student groups on the same set of GPUs in the node, which run in parallel to process different individual samples respectively and concurrently. Specifically, if the size of the student group is $S$ and total GPU number is $G$, it always assign the $i$-th student of $j$-th student group on the specific GPU as follows:
(5) $\text{GPU}_{id}=(i+j\times S)\%G.$
Therefore, if $i+j\times S>G$, the i-th student in the j-th student group is co-located on the same GPU with other students in the previous groups. Student Parallelism employs NVIDIA Multi-Process Service (MPS) (mps, ) to make multiple students from different groups share the same GPU without interference. It can guarantee different students on the same GPU have separate computation units and memory bandwidth to run concurrently.
Param# SST-2 RTE MNLI CoLA(MMS) QQP MRPC(F1) STS-B(PCC) QNLI Overall Score
*BERT-base-12L768D 109 M 92.7 67 83.6 52.8 89.6 88.6 89 91.8 *81.89(100%)
I-BERT-12L768D 109 M 91.6 65.7 81.3 49.1 87.1 85.2 86.9 87.4 *79.29(96.82%)
DeeBERT-12L768D 109 M 91.5 66.1 80 43.4 87.3 85.2 - 87.9 77.34(94.45%)
DistillBERT-6L768D 67 M 91.3 58.4 81.1 49 88.1 86.9 86.9 88.9 78.83(96.26%)
DynaBERT-6L192D 9 M 92 63.2 82.15 43.7 90.4 81.4 87 88.5 78.54(95.92%)
TinyBERT-4L312D 15 M 87.8 60.8 76.9 44.1 87.7 85.8 83.3 86 76.55(93.48%)
PKD-BERT-3L768D 46 M 87.5 58.2 76.3 24.8 87.8 80.7 79.8 84.7 72.48(88.51%)
BERT-2L256D 10 M 87.1 54.6 74.7 23.2 87 80.3 86 84.4 72.16(88.12%)
Academus-2L (pruned to 1) 10 M 91.3 64.9 77.3 42.3 88.1 86.7 85.2 88.2 78.00(95.25%)
Academus-2L (best pruned) 10*m M 91.8 65.9 78.2 43.2 88.7 86.9 86.4 89.8 78.86(96.31%)
*BERT-large-24L1024D 335 M 93.2 70.4 86.6 60.6 91.3 89.3 90 92.3 83.06(100%)
Academus-6L768D (best pruned) 67*m M 92.1 61.4 82.5 51.2 89.2 88.9 87.2 89.7 80.28(96.65%)
*DeBERTa-xlarge-48L1024D 658 M 97 93.1 91.3 70 92.3 94.3 92.8 95.1 90.40(100%)
Cocktail $\sum p_{i}$ 94.2 72.1 87.1 62.1 91.5 89.7 - 93.4 84.30(93.25%)
Academus-12L768D (best pruned) 109*m M 94.1 78.7 88.8 59.3 91.9 90.1 91.5 93.1 85.94(95.06%)
Table 1. Measurement and comparison on prediction quality: * means the large model with full depth and width; xLxxxD means the model has x layers and xxx dimensions in the hidden state; the bold results means the better overall score than the other baselines without full depth or width; and the underline means the target quality achieved.
During the inference, resource management keeps tracking the intensity of dynamic online workloads(i.e., the request per second) to adjust the student number accordingly. It temporarily decrease the student number (§ 3.2.2) to run more student groups for workload burst, instead of reserving some idle GPUs in advance. And after workload bursts, Student Parallelism can increase the student number to fully utilize all GPUs for high accuracy. In the initialization, resource management conduct profiling to determine the proper hyper-parameters of student allocation, e.g., the optimal student number sharing the same GPU (more details in the Appendix).
3.3.2. Inference Sample Dispatching
To reduce the extra costs of idle waiting and paddings for batching from online workloads, Student Parallelism enable the immediate and concurrent inference on different individual samples. If there is any idle student group and no buffered samples, the dispatcher can immediately send out the arrived sample for inference without waiting for batching. Since different individual samples do not arrive at the same time, different student groups can concurrently process them by overlapping between data transferring to GPU and inference computation as shown in Figure 5. Only if there are massive concurrent samples occupying all student groups, the length-aware buffer temporarily stores the new samples, which merges the samples in the same length bin into small data batches with a few paddings. To efficiently merge samples on the fly, the length-aware buffer combines the queue with a dictionary whose key-value pair is length bin and queue position respectively (details in the Appendix).
4. Evaluation
In this section, we first introduce the experiment settings. Then we evaluate the overall performances of Student Parallelism. Finally, we conduct ablation experiments to study the influences of model hyperparameters and the sources of the improvement. From the evaluation, we can draw the following important conclusions:
  • Student Parallelism can effectively reduce more model depth, while maintaining the best accuracy among small models without full depth or width. And adaptive student pruning can achieve comparable accuracy with even only one student pruned from all students.
[…]
  • Based on adaptive student pruning, it can improve throughput up to 22.27×22.27$\times$22.27 × compared to different methods.
Method Model Config.(Param#) Avg. Latency 95% Tail Latency Throughput per GPU
Value (ms) Reduction (%) Value (ms) Reduction (%) Value(#/s) Improvement ($\times$)
*BERT-base 12L768D (109 M) 11.6 0.00% 11.7 0.00% 691.5 1.00
I-BERT 12L768D (109 M) 4.9 -57.8% 9.5 -18.80% 2115.99 3.06
DeeBERT 12L768D (109 M) 6.8 -41.38% 11.5 -1.71% 941.4 1.36
DynaBERT 6L192D (9 M) 8.3 -28.45% 9.6 -17.95% 7538.7 10.9
DistillBERT 6L768D (67 M) 8.4 -27.59% 9.7 -17.09% 1508.1 2.18
TinyBERT 4L312D (15 M) 7.3 -37.07% 9.7 -17.09% 5873.6 8.49
PKD-BERT 3L768D (46 M) 6.8 -41.38% 9.5 -18.80% 2781.3 4.02
BERT-2L 2L256D (10 M) 4.5 -61.21% 9.4 -19.66% 14420.3 20.85
Academus-2L1S 2L256D1S (10 M) 2.8 -75.86% 3.6 -69.23% 15398.6 22.27
Academus-2L3S 2L256D3S (30 M) 3.0 -74.13 3.7 -68.37% 520191 7.52
*BERT-Large 24L1024D (335 M) 21.1 0.00% 21.3 0.00% 239.6 1.00
Academus-6L1S 6L768D1S (67 M) 6.3 -70.14% 7.1 -66.67% 1508.1.5 6.29
Academus-6L3S 6L768D(1-3)S (201 M) 6.6 -68.87% 7.3 -65.72% 512.4 2.14
*DeBERTa-xlarge 48L1024D (658 M) 40.5 0.00% 41.4 0.00% 160.9 1.00
Cocktail - 21.4 -47.16% 22.3 -46.14% 240.4 1.49
Academus-12L1S 12L1024D1S (109 M) 11.8 -70.86% 12.6 -69.57% 621.5 3.86
Academus-12L3S 12L1024D3S (327 M) 12.1 -70.12% 12.8 -69.08% 226.5 1.41
Table 2. Measurement on inference efficiency: * means the large model with full depth and width ; xLxxxD means the model has x layers and xxx hidden dimensions, and xS means the student number is x; the bold results means the better efficiency performance than the others without full depth or width; italic indicates the method fails to satisfy the target accuracy.
4.1. Experiment Settings
4.1.1. Implementation Detials
In training, we consider all 24 compact BERT variants (turc2019well, ) and all DEBERTA variants as the student candidates to find the shallowest one for different teachers to reach target accuracy. In practice, we use the BERT-2L256D (turc2019well, ) for BERT-base, BERT-6L768D for BERT-large, and Deberta-12L768D for DEBERTA-xlarge as the backbone network and initialization of student models respectively. Following previous works (jiao2019tinybert, ; DBLP:conf/acl/XinTLYL20, ), we conduct the same data augmentation on the down-stream datasets. We conduct a grid search on various training hyperparameters to find the best for every dataset, including epoch number, batch size, learning rate, $\lambda$ weights, and data sampling ratio. And in practice, we optimize the $\alpha_{i}$ together with the student, since they are both differentiable. We run inference on 4 servers each with 4 Geforce 3090 GPUs and more implementation details about online inference systems can be found in Appendix.
4.1.2. Datasets and Metrics
We follow previous works (DBLP:conf/acl/XinTLYL20, ; sanh2019distilbert, ; hou2020dynabert, ) to use the famous GLUE benchmark (wang2018glue, ). It consists of 8 text datasets in various text mining (SST-2, CoLA, MNLI, and RTE) and information retrieval (MRPC, QQP, STS-B, and QNLI) tasks. We report various prediction quality metrics recommended by GLUE on the validation set of different tasks, including accuracy, F1 score, Matthews Correlation Coefficient (MCC) and Pearson correlation coefficient (PCC). Additionally, we also build our workload generator upon the real production trace of twitter (khaleq2018cloud, ) to feed the inference system. For discriminative text mining, we report the average latency and the 95% tail latency. And we report throughput for information retrieval tasks to better show how long users will wait for the ranked results.
4.1.3. Baselines
We compare our Student Parallelism with various types of baselines related to BERT-like models. First, we consider the finetuned 12-layer BERT-base-12L768D (devlin2018bert, ), 24-layer BERT-large-24L1024D (devlin2018bert, ), and 48-layer DeBERTa-xlarge-48L1024D (he2021deberta, ) as the original teachers. Then we conduct quantization to get I-BERT-12L768D with the same model architecture. We train the 2-layer, 6-layer, and 12-layer versions of Student Parallelism for these teachers respectively, reducing over 75% model depth. And we make the DeeBERT-12L768D (DBLP:conf/acl/XinTLYL20, ) the baseline for the adaptive computing. We use the 3-layer PKD-BERT-3L768D (sun2019patient, ), 4-layer TinyBERT-4L312D (jiao2019tinybert, ), and 6-layer DistillBERT-6L768D (sanh2019distilbert, ) as the representatives of knowledge distillation. For pruning, we choose the fastest version of DynaBERT-6L192D (hou2020dynabert, ) satisfying the target prediction. We also bring in extremely shallow finetuned models like BERT-2L256D (turc2019well, ) as the vanilla 2-layer model. However, all these baseline works can only optimize finetuned 12-layer BERT-base model. We further compare our work with Cocktail based on bagging ensemble of all the above-mentioned models except our Student Parallelism to approach the performance of 48-layer DeBERTa-xlarge. By following the previous works (zhang2019mark, ; gunasekaran2022cocktail, ), all baselines employ the dynamical batch size and ragged batch to realize adaptive batching for practical online workloads. We set the overall time budget as 10ms to dynamically set the batch size and waiting window according to profilings and online workloads (zhang2019mark, ).
4.2. Overall Performance
In this subsection, we first verify if Student Parallelism can maintain the accuracy. Then we show its advantages in inference efficiency, including average latency, tail latency and throughput.
4.2.1. Prediction Quality
As shown in Table 1, Student Parallelism can achieve the best overall score up to 96.32% of BERT-base with only two layers, and it is the only one compressing BERT-large and DeBERTa-xlarge into their one-quarter model depth while reaching the target accuracy. The results show other baselines can only reduce about half the layers of the BERT-base to maintain the accuracy, which reach the target score (95% overall score of the teachers, as recommended in MLPerf inference benchmark (reddi2020mlperf, )). Compared with tiny baselines with less than 6 layers, Student Parallelism has significant advantages in the prediction qualities. Especially, the 3-layer PKD-BERT and 2-layer BERT-2L512D can only maintain less than 90% overall score. Student Parallelism is also better than any 6-layer or deeper reduced baselines with over 3$\times$ larger model depth. It is even comparable with I-BERT always having full depth and width. Because of improved generalization from adaptive student pruning, our work can use even one student to handle workload burst and reach a competitive overall score (95.25% of the BERT-base) compared with 6-layer baselines. Additionally, if using deeper student models, Student Parallelism also can reduce the model depth of larger BERT-large and DeBERTa-xlarge by 75%, while the others fail. And Cocktail cannot reach the target of DeBERTa-xlarge even with 24-layer BERT-large. Since DeeBert and Cocktail rely on the classification probability to work, we omit their results in the regression dataset STS-B.
4.2.2. Latency
[…]
4.2.3. Throughput
In information retrieval tasks, like the search engine and QA system, the user has to wait to get the final result until dozens of document candidates are scored for the relevance of the query. Hence, we report the throughput per GPU for best-pruned students and the single student to compare with other baselines on the same number of GPUs. Because some models have been out of memory, we set the batch size as 64 for all baselines. The last column of Table 2 records the overall throughput / GPU number for all methods in the MRPC dataset. Our work can improve throughput over the others by up to 22.27×22.27$\times$22.27 ×, because it can leverage adaptive student pruning to use only one small student whose generalization has been improved (§ 4.2.1). Meanwhile, even the best pruned of students (3 in total) can achieve comparable throughput with some competitive baselines like TinyBERT for better accuracy if the workload is not so heavy. Except for the embedding layer, all model parameters mean computations, so the throughput negatively correlates with the model size. Even with fast integer calculation, the full model size of I-BERT limits its throughput improvement to 3×3$\times$3 ×. The 6-layer DynaBERT is the most competitive one of the small baselines reaching the target accuracy, because it reduces model width to be $25\%$. The vanilla BERT-2L256D has high throughout, but it is far away from the target accuracy. And Cocktail suffers from low throughput per GPU since the bagging ensemble has to use multiple GPUs.
4.3. Abalation Studies
In this section, we further conduct ablation studies in Student Parallelism, including hyperparameters tuning, accuracy analysis experiment, and latency analysis experiment.
4.3.1. Model Hyperparameter Tuning
(a) Student Number
(b) Model Depth
Figure 6. Hyperparameter influences on prediction quality: (a) Different student number comparison. (b) Different layer number influences comparison
Since student number and model depth are key hyperparameters influencing the prediction quality, we conduct experiments to measure how they affect the prediction quality. According to the results, we verify that adaptive student pruning can mprove the generalization of the remained students and reduce the student number with minimal accuracy loss. Furthermore, we can even further improve prediction quality with larger model depth with the relaxed latency requirement.
MRPC (F1)
Student Parallelism 86.9
w/o pruning 86.8
w/o stacking distillation 86.0
replacing boosting with bagging 85.6
Table 3. Prediction quality ablation experiment results
Figure 6(a) shows that adaptive student pruning consistently outperforms the direct boosting ensemble of the same number of students without pruning in overall GLUE scores. The results verify that adaptive student pruning can effectively improve the generalization as a pruning technology. Because it improves the generalization of the remained student, 3 students are the best pruned to get the highest test performance in the most datasets. And even only one student pruned from all 6 students can maintain a relatively high score. However, when we directly train the boosting ensemble of $m$ students without pruning, the limited total student number has more difficulties in finding the optimal solution.
As shown in Figure 6(b), our Student Parallelism can further improve the prediction with a deeper student when learning from BERT-base in all datasets. The 4-layer students distilled from BERT-base can significantly improve the prediction results. Student Parallelism can even get a similar F1 score to the original 12-layer BERT-base if using 6-layer students. And compared with other best-optimized baselines that have the same layers (such as BERT-2L256D, TinyBERT, and DistillBERT), our Student Parallelism can consistently outperform them in overall GLUE scores, while having the same layer number and similar inference latency.
4.3.2. Accuracy Analysis
As shown in Table 4, we first conduct accuracy ablation experiments to verify the effectiveness of our training methods to maintain accuracy on the MRPC dataset, as stated in Table 3. When we directly use all trained students without adaptive student pruning, we observe that F1 score decreases a little from 86.9 to 86.8. After removing the stacking distillation for better refinement, the F1 score further drops to 86.0, which is slightly lower than the TinyBERT-4L312D. We further replace the boosting ensemble with a simple bagging ensemble of multiple small students, and the performance will further decrease. The same trend also happens on the other datasets.
4.3.3. Latency Analysis
We also make breaking-down experiments to evaluate the influences and contributions of different design parts on inference latency. The results show that latency increases dramatically when we increase the student’s model depth to 4 layers, which indicates that reducing serial computation is the key to low latency. After we pad all data samples into the maximum length, the average latency increase to be similar to tail latency. It means that the variances of Student Parallelism latency mainly come from the irregular sentence lengths. We further employ the waiting queue to batch enough samples for data parallelism, then average latency increase a lot and tail latency even get close to the overall time budget.
avg (ms) tail (ms)
Student Parallelism 2.8 3.6
4-layer students 4.5 5.2
with padding 5.1 5.3
with waiting queue 7.5 9.5
Table 4. Latency ablation experiment results
5. Conclusion
We propose Student Parallelism to adopt student parallelism by distilling the deep BERT-like model into an equivalent group of parallel shallow students. Every student model can have decoupled and homogenous architecture of fewer layers for low inference latency, while all student models work together to maintain accuracy via boosting ensemble and stacking distillation. And adaptive student pruning can temporally drop some students to improve the throughput for workload burst. We further make specialized system designs for student parallelism to minimize waiting and padding. Comprehensive experiments have verified the effectiveness of our work in terms of latency, throughput, and accuracy.
Appendix A Convergence Analysis
In this section, we conduct the convergence analysis on our training methods based on boosting ensemble, which shows the residual training loss converges to the infimum as more students are added during the training. To prove the convergence, we can first prove our sequential training of students is one special case of Anyboost [25, 26], then we can reuse the convergence properties and proofs of Anyboost in our proposed Student Parallelism.
Specifically, we take a non-parametric approach to view our sequential training of students as the numerical optimization problem in function space. And we consider $B(x))$ evaluate at each point $x$ as a ”parameter” and seek to minimize the following MSE error:
(6) $L(T(x),B(x)))=\frac{1}{2}|T(x)-B(x))|^{2},$
to make the ensemble of students generates a similar final representation with the original teacher $T(x)$. In function space, there are an infinite number of such parameters, but in data sets only a finite number $\{B(x)_{i})\}_{1}^{N}$ are involved. Following the numerical optimization paradigm we take the solution to be
(7) $F^{*}(x)=\sum_{i=0}^{M}f^{(i)}(x),$
where $f^{(0)}(x)$ is the initial student $S^{(0)}(x)$ directly mimics the teacher, and any $f^{(i!=0)}(x)$ is the incremental function (”step” or ”boost”) defined by the optimization method as follows:
(8) $f^{(i)}(x)=-\alpha_{i}g^{(i)}(x).$
And $g^{(i)}(x)$ is the gradient decent direction of $L(\cdot)$, namely:
(9) $\begin{split}g^{(i)}(x)=T(x)-B^{(i-1)}(x)\end{split}$
Therefore, our main loss $L_{\text{boost}}$ just makes the added i-th student $S^{(i)}(x)$ fit the $g_{i}(x)$:
(10) $\begin{split}L_{\text{boost}}(x)=\frac{1}{2}\left\|T(x)-B^{(i-1)}(x)-S^{(i)}(x)\right\|^{2}=\frac{1}{2}\left\|g^{(i)}(x)-S^{(i)}(x)\right\|^{2}\end{split}$
Although we apply other regularization like stacking distillation, the final representation of the student is only related with the above loss. Since the added i-th student $S^{(i)}(x)$ is trained to fit $g^{(i)}(x)$ in the final output and $\alpha_{i}$ is determined by the line serach, the sequential training of students in Student Parallelism belongs to Anyboost [25, 26]. As a result, the boosting ensemble of students to mimic the teacher in Student Parallelism should share the same properties with Anyboost to have the following convergence theorem:
Theorem A.1.
The MSE in Student Parallelism has the lower bound 0, and it is Lipschitz differentiable loss function (for any L ¿ 1, we always have $|\bigtriangledown L_{\text{boost}}(x_{1})-\bigtriangledown L_{\text{boost}}(x_{1})|<L|x_{1}-x_{2}|$). Let $F^{(0)},F^{(1)},...$ be the sequence of combined hypotheses generated by the Student Parallelism training algorithm, using small enough step-sizes $\alpha_{i}:=-\frac{\langle\bigtriangledown L_{\text{boost}}(T,B^{(i-1)}),S^{(i)}\rangle}{L|S^{(i)}|^{2}}$. Then our sequential training of students either halts on round T with $-\langle\bigtriangledown L_{\text{boost}}(T,F^{(i-1)}),S^{(i)}\rangle\leq 0$, or $L_{\text{boost}}(F^{(i)})$ converges to some finite value $L^{*}$, in which case
(11) $\lim_{i}\langle\bigtriangledown L_{\text{boost}}(T,B^{(t)}),S^{(t+1)}\rangle=0.$
Because the detailed proof can be found in the previous theory work [25, 26], we skip the proof for the convergence theorem.
Appendix B Implementation Details
B.1. Testbed and Environment
We use 4 servers each with 4 NVIDIA Geforce 3090 GPUs, Intel Xeon(R) E5-268340 CPU, and 96 GB memory for all evaluations. The server is supported by a Network File System (NFS) to store the models and datasets in the storage node. And it runs on the Ubuntu 18.04 operating system. Except the global dispatcher, every server has its own logical controller and computation workers. And we implement the Student Parallelism with Python. Specifically, we use the PyTorch [29] as our DL framework, HuggingFace Transformers [42] for the model implementations, , NCCL [15] for the communications among students.
B.2. Offline Training Implementation
In the offline training, we consider all 24 compact BERT variants [39] and all DEBERTA variants as the student candidates to find the lowest one for different teachers. In practice, we use the BERT-2L256D [39] for BERT-base, BERT-6L768D for BERT-large, and Deberta-12L768D for DEBERTA-xlarge as the backbone network and initialization of student models respectively. Following previous works [16, 43], we conduct the same data augmentation on the down-stream datasets. We conduct a grid search on various training hyperparameters to find the best for every dataset, including epoch number, batch size, learning rate, $\lambda$ weights, and data sampling ratio. And in practice, we optimize the $\alpha_{i}$ together with the student, since they are both differentiable.
B.3. Online Inference Implementation
For online inference, the global dispatcher to assign samples among different nodes in the round-robin method. And every node employs message queues to dispatch samples and collect the results in the master-worker architecture. The master process is responsible for logical controlling, including resource management, the length-aware buffer, and task dispatching. And workers host all concurrent student groups on all the GPUs. Once a user submits an inference request, the master preprocesses, buffers, and then dispatches the inference samples. And any idle worker can fetch the inference task from the message queue and return the results. We employ the Multi-Process Service (MPS) [28] to make multiple students of different groups share the same GPUs without interference.
B.3.1. Hyperparameters for Student Allocation
The resource management determines the proper hyperparameters of student allocation. In the initialization, it conducts profiling and grid searching to find the optimal student number sharing the same GPU and the maximum batch size of each student group (used in § B.3.2). Both hyperparameters are set to maximize the total throughput without increasing latency due to lacking computation capability. During the inference, it dynamically adjusts the student number to host different numbers of student groups according to the workload intensity and resource usage. It keeps tracking the buffer size (decribled in § B.3.2) and the idle group number as indicators of workload intensity and resource usage. If the buffer size reaches the threshold (i.e. the total student group number), Student Parallelism drops one last-trained student models to start more groups, so that the waiting time of buffered samples will not exceed the inference time. And if the buffer is empty and the total student number of all idle groups is larger than occupied group number for a certain time window (e.g., 2 minutes), it means there are enough free resources to increase the student number by one for all the occupied student groups. Then Student Parallelism can make better use of idle computation resources to improve accuracy.
B.3.2. Dispatch with Length-Aware Buffer
Figure 7. Length-aware Buffer: New sequence sample (e.g., the red one whose length is 30) can be merged with the previous element into a small batch, since they share the same bin and the element size is not full. Otherwise, the new sequence sample (e.g., the blue one whose length is 120) is added to the buffer tail as a new element.
Although student parallelism reduces the necessity of data parallelism via direct inference on individual samples, we further propose the length-aware buffer to merge the concurrent samples of similar lengths into small batches, in case that all student groups are occupied. Different from the waiting queue that keeps waiting until enough samples arrive, the length-aware buffer only temporarily stores the new samples if all student groups have been occupied by previous samples. If idle student group is available and the length-aware buffer is empty, Student Parallelism directly sends out the new sample for the direct inference. Once the inference sample arrives, we check which length bin it belongs to and pad or clip it into the maximum length of the bin. Only if the buffer is not empty, it tries to merge the new sample and others in the same length bin into a small data batch as one buffer element on the fly. Whenever any running group is idle, the dispatcher immediately sends out the first buffer element (a sample or small batch) for inference without any delay. We set the maximum buffer size as the total number of current student groups, so the range of buffer time is from zero to the inference latency.
Figure 7 shows the details about how the length-aware buffer efficiently generates small batches of samples in similar lengths on the fly. We equally split the full length range (i.e., 0-128) into 16 bins at every step of 8 tokens. Then we set the bins and pointer as the key-value pair in the dictionary to quickly index samples of different length bins. When a new sample arrives, we check the dictionary to see if any sample belongs to the same length bin in the buffer. If so and the element size is not the maximum (e.g., 4, which is also determined by profiling), we can get the element pointer from the dictionary to merge them as one buffer element. Otherwise, we append the new sample as the tail element in the buffer and update its length bin and pointer in the dictionary. When the first buffer element is dispatched, we delete its key-value pair in the dictionary. Obviously, all the operators of the length-aware buffer only has $O(1)$ time complexity due to the hash map.